# deCIFer: Crystal Structure Prediction from Powder Diffraction Data using Autoregressive Language Models

### v2

@@ Appendix A Appendix
The code for training and using the deCIFer model is open source and released under the MIT License. Source code is available here: https://github.com/FrederikLizakJohansen/deCIFer.
@@
To enhance the transformer model to process CIFs effectively, we standardized all CIFs in the dataset. Inspired by CrystaLLM Antunes et al. [2024], we employed similar pre-processing and tokenization strategies, incorporating additional steps to ensure that CHILI-100K Friis-Jensen et al. [2024] was aligned to the standardized format of NOMA, by the removal certain details such as oxidation states and partial occupancies. We employ the following steps:
@@
    Uniform Structure Conversion: CIFs were converted to pymatgen.Structure Ong et al. [2013] objects to provide a consistent base representation.
@@
    Standardized CIF Regeneration: Using pymatgen.CifWriter Ong et al. [2013], CIFs were regenerated to ensure uniform formatting, eliminate customs headers, etc.
@@
    Oxidation State and Occupancy Filtering: Oxidation state refers to the charge of an atom within a compound, which can vary depending on chemical bonding. Occupancy indicates the fraction of a particular atomic site that is occupied in the crystal structure (e.g., a value of 1.0 represents a fully occupied site, while 0.5 indicates partial occupancy). All traces of oxidation states were removed, and only crystal structures with full occupancy were retained. This ensures consistency by aligning CHILI-100K Friis-Jensen et al. [2024] with the standardized format of NOMA Antunes et al. [2024].
@@
Figure 9: Illustration of a CIF after applying the pre-processing and standardization steps described. Also shown are the corresponding unit cell representation using VESTA Momma and Izumi [2008] for visualization and the simulated PXRD profile (with $\sigma^{2}=0.00$ and FWHM=0.01). The red highlight in the CIF indicates where the original symmetry operators were replaced during pre-processing and would be restored for evaluation.
@@
Following Section 2.1, we start with the discrete diffraction peak data: ${\mathcal{P}}={(q_{k},i_{k})_{k=1}^{n}}$, where each $q_{k}$ is the center of a reflection peak, and $i_{k}$ is the associated peak intensity. To simulate experimental effects, we apply transformation $\tau\sim{\mathcal{T}}$, which includes peak broadening and additive noise.
@@
In practice, we use the XRDCalculator from the pymatgen library Ong et al. [2013] for generating the initial discrete peak data ${\mathcal{P}}$. For training, we sample $Q$-values in $[Q_{\mathrm{min}},Q_{\mathrm{max}}]$ at increments of $Q_{\mathrm{step}}$. We then apply random transformations $\tau$ during model training. Specific parameters for FWHM and $\sigma_{\mathrm{noise}}$ are listed in Table 5.
@@
  • We use a Voronoi-based nearest-neighbour algorithm implemented in the CrystalNN module of pymatgen Ong et al. [2013] to identify bonded atoms.
@@
  • Analyzing the space group symmetry using the SpacegroupAnalyzer class in pymatgen Ong et al. [2013], which employs the spglib Togo and Tanaka [2018] library.
@@ A CIF file is deemed valid if all the above checks are satisfied:
The Match Rate (MR) quantifies how many generated structures successfully match their corresponding reference structures, as determined by StructureMatcher from the pymatgen library Ong et al. [2013]. Two structures are considered a match if their compositions, lattice parameters, atomic coordinates, and symmetry are sufficiently similar, according to the tolerances set in StructureMatcher. For the implementation of deCIFer, we follow the example set by CrystaLLM Antunes et al. [2024], using the parameter values:
@@
Figure 11 illustrates the NOMA dataset. Figure 12 illustrates the statistics of the curated CHILI-100K Friis-Jensen et al. [2024] dataset.
Figure 11: Statistical overview of the NOMA Antunes et al. [2024] dataset (2,283,346 total samples), showing the distribution of space group frequencies, the number of elements per unit cell, elemental occurrences and CIF token lengths (indicating the percentage of CIFs with larger token sequences than the context length of 3076)
Figure 12: Statistical overview of the curated CHILI-100K Friis-Jensen et al. [2024] dataset (8201 total samples), showing the distribution of space group frequencies, the number of elements per unit cell, elemental occurrences, and CIF token lengths (indicating the percentage of CIFs with larger token sequences than the context length of 3076).
@@
Figures 15, 16, and 17 provide additional insights into deCIFer’s performance. Table 6 shows a detailed breakdown of validity metrics for the NOMA test set and CHILI-100K test set evaluated on two in-distribution (ID) scenarios and one out-of-distribution (OOD) scenario for the PXRD input.
Figure 15: Distribution of $R_{wp}$ and WD for deCIFer on the NOMA test set, presented as violin plots with overlain boxplots; the median is shown for each distribution. Presented are four in-distribution transformations of the input PXRD profiles and three out-of-distribution transformations.

### v3

@@ Appendix A Appendix
The code for training and using the deCIFer model is open source and released under the MIT License. Official source code is available here: https://github.com/FrederikLizakJohansen/deCIFer.
@@
To enhance the transformer model to process CIFs effectively, we standardized all CIFs in the dataset. Inspired by CrystaLLM [Antunes et al., 2024], we employed similar pre-processing and tokenization strategies, incorporating additional steps to ensure that CHILI-100K [Friis-Jensen et al., 2024] was aligned to the standardized format of NOMA, by the removal certain details such as oxidation states and partial occupancies. We employ the following steps:
@@
    Uniform Structure Conversion: CIFs were converted to pymatgen.Structure [Ong et al., 2013] objects to provide a consistent base representation.
@@
    Standardized CIF Regeneration: Using pymatgen.CifWriter [Ong et al., 2013], CIFs were regenerated to ensure uniform formatting, eliminate customs headers, etc.
@@
    Oxidation State and Occupancy Filtering: Oxidation state refers to the charge of an atom within a compound, which can vary depending on chemical bonding. Occupancy indicates the fraction of a particular atomic site that is occupied in the crystal structure (e.g., a value of 1.0 represents a fully occupied site, while 0.5 indicates partial occupancy). All traces of oxidation states were removed, and only crystal structures with full occupancy were retained. This ensures consistency by aligning CHILI-100K [Friis-Jensen et al., 2024] with the standardized format of NOMA [Antunes et al., 2024].
@@
Figure 7: Illustration of a CIF after applying the pre-processing and standardization steps described. Also shown are the corresponding unit cell representation using VESTA [Momma and Izumi, 2008] for visualization and the simulated PXRD profile (with $\sigma^{2}=0.00$ and FWHM=0.01). The red highlight in the CIF indicates where the original symmetry operators were replaced during pre-processing and would be restored for evaluation.
@@
Section A.4 Attention Masking Strategy
Figure 8 provides a detailed visualization of the attention masking strategy employed in our model. It illustrates the log-mean attention weights (averaged over all heads) for a sample sequence, highlighting the isolation of CIFs through attention masking. The figure also demonstrates how the embeddings of the second CIF attend to the conditioning PXRD embedding. Lighter shades in the figure correspond to stronger attention values.
@@
Following the methods section, we start with the discrete diffraction peak data: ${\mathcal{P}}={(q_{k},i_{k})_{k=1}^{n}}$, where each $q_{k}$ is the center of a reflection peak, and $i_{k}$ is the associated peak intensity. To simulate experimental effects, we apply transformation $\tau\sim{\mathcal{T}}$, which includes peak broadening and additive noise.
@@
In practice, we use the XRDCalculator from the pymatgen library [Ong et al., 2013] for generating the initial discrete peak data ${\mathcal{P}}$. For training, we sample $Q$-values in $[Q_{\mathrm{min}},Q_{\mathrm{max}}]$ at increments of $Q_{\mathrm{step}}$. We then apply random transformations $\tau$ during model training. Specific parameters for FWHM and $\sigma_{\mathrm{noise}}$ are listed in Table 4.
@@
  • We use a Voronoi-based nearest-neighbour algorithm implemented in the CrystalNN module of pymatgen [Ong et al., 2013] to identify bonded atoms.
@@
  • Analyzing the space group symmetry using the SpacegroupAnalyzer class in pymatgen [Ong et al., 2013], which employs the spglib [Togo and Tanaka, 2018] library.
@@ A CIF file is deemed valid if all the above checks are satisfied:
The Match Rate (MR) quantifies how many generated structures successfully match their corresponding reference structures, as determined by StructureMatcher from the pymatgen library [Ong et al., 2013]. Two structures are considered a match if their compositions, lattice parameters, atomic coordinates, and symmetry are sufficiently similar, according to the tolerances set in StructureMatcher. For the implementation of deCIFer, we follow the example set by CrystaLLM [Antunes et al., 2024], using the parameter values:
@@
Figure 10 illustrates the NOMA dataset. Figure 11 illustrates the statistics of the curated CHILI-100K [Friis-Jensen et al., 2024] dataset.
Figure 10: Statistical overview of the NOMA [Antunes et al., 2024] dataset (2,283,346 total samples), showing the distribution of space group frequencies, the number of elements per unit cell, elemental occurrences and CIF token lengths (indicating the percentage of CIFs with larger token sequences than the context length of 3076)
Figure 11: Statistical overview of the curated CHILI-100K [Friis-Jensen et al., 2024] dataset (8201 total samples), showing the distribution of space group frequencies, the number of elements per unit cell, elemental occurrences, and CIF token lengths (indicating the percentage of CIFs with larger token sequences than the context length of 3076).
@@
Data-stratification
We extract the space group number from each CIF (ranging from $1$ to $230$) and group these into bins of size ten (e.g., 111110101010, 1111111120202020, etc.). This heuristic aims to preserve the overall symmetry distribution across splits while reducing the risk of data leakage from structurally similar entries appearing in multiple subsets. While this does ensures coverage across symmetry classes and even representation of crystal systems across the splits, it does not reflect the most intuitive or principled grouping scheme. In particular, it does not account for finer-grained biases that may be embedded in crystal symmetry or composition distributions. This points to a broader issue in materials datasets: statistical artefacts, such as the "Rule of Four" or symmetry clustering [Gazzarrini et al., 2024], can introduce shortcuts that models may learn, reducing generalisation and interpretability [Palgrave, 2024]. Future work should explore alternative stratification strategies (e.g., stratified sampling based on structural descriptors) to better assess generalisation.
@@
Section A.11 Baseline Comparison with and without PXRD
With regards to the baseline comparison in Table 1, deCIFer is explicitly conditioned on PXRD data, while the baseline models are conditioned only on composition or latent priors. This distinction means that the comparisons are not direct but instead reveal the relative value of PXRD conditioning. In particular, PXRD conditioning can improve structure prediction when the diffraction signal is rich in structural information, but may introduce ambiguity or conflict with the model’s learned priors in cases where the PXRD pattern is noisy or minimally informative. Perov-5 and Carbon-24 provide strong tests of PXRD conditioning due to their polyhedral complexity and carbon-based structural diversity, where diffraction features can directly inform the model. In contrast, MP-20 is drawn from the Materials Project, a dataset where composition-only models may benefit from learned priors due to the high representation of standard chemistries. MPTS-52 further challenges models with low-symmetry structures, where the PXRD signal can become ambiguous, making it difficult for a model to resolve atomic positions purely from the diffraction pattern.
@@
Figures 16, 14, and 15 provide additional insights into deCIFer’s performance. Table 5 shows a detailsed breakdown of the validity metrics for the NOMA test set corresponding to the results in Figure 3. Table 6 shows a detailed breakdown of validity metrics for the NOMA test set and CHILI-100K test set evaluated on two in-distribution (ID) scenarios and one out-of-distribution (OOD) scenario for the PXRD input.
Section A.13 Future Work
Another direction could be to integrate reinforcement learning from human feedback (RLHF) to guide the model more directly toward generating accurate and chemically valid structures [Ziegler et al., 2019]. By defining a reward function tailored to properties such as low $R_{\mathrm{wp}}$ values, structural integrity, and adherence to chemical constraints, and interaction with a human expert, RLHF could further refine the model’s outputs.
A complementary direction for future improvement lies in increasing the diversity of the training data. While NOMA offers excellent scale, its distribution is skewed toward high-symmetry and well-sampled structures, which limits model generalisation to more uncommon systems. CHILI-100K, with its broader structural complexity and greater representation of low-symmetry crystals, could serve as a valuable training supplement. Mixing synthetic and experimental data, or applying curriculum learning strategies that gradually expose the model to under-represented symmetry groups, could improve generalisation and robustness, particularly for monoclinic and triclinic systems, which remain challenging as seen in Figure 4.
none U-deCIFer 99.82 98.87 94.30 99.47 93.49 0.00
deCIFer 99.42 98.85 93.69 99.46 92.66 5.01
comp. U-deCIFer 99.87 99.09 94.40 99.46 93.78 49.30
deCIFer 99.68 99.21 94.37 99.55 93.73 91.50
comp.+s.g. U-deCIFer 99.85 98.88 94.51 99.47 93.72 87.07
deCIFer 99.74 99.26 94.38 99.58 93.90 94.53
@@
Figure 16: Distribution of $R_{wp}$ for deCIFer on the NOMA- and CHILI-100K test set, presented as violin plots with overlain boxplots; the median is shown for each distribution. Presented are four in-distribution transformations of the input PXRD profiles and three out-of-distribution transformations.